# Online Fair Allocations with Binary Valuations and Beyond

### v1

@@ 4.2.1 Additive Binary Costs
Output: An complete EF1 allocation $A$
1 Initialize: let $\pi=(\pi_{1},\cdots,\pi_{n})$ be an arbitrary order of $n$ agents;
2Let $A=(\emptyset,\ldots,\emptyset)$;
@@ 14      $\pi=(\pi_{2},\cdots,\pi_{i-1},\pi_{i+1},\cdots,\pi_{n},\pi_{1})$;
Given an online fair allocation instance $I=(T,N,(c_{i})_{i\in N})$, where $c_{i}$ is additive and binary for all $i\in N$. Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ be the set of items that have already arrived, denoted by ${\bf A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $k$. For every item $e\in T$, if there exists one agent such that the cost of $e$ to some agent is 00, the algorithm must allocate $e$ to some agent $a$ with $v_{a}(e)=0$. Otherwise, the costs of all agents are $1$. Regardless of how it is allocated, the increment brought to utilitarian social welfare is $1$, so $A^{k}$ is a min-USC allocation.
@@ 4.2.2 Supermodular Binary Costs
For supermodular binary cost functions, where for any agent $i\in N$, $c_{i}$ is supermodular and the marginal cost of any item is 0 or 1, we first obtain a significant negative result regarding the approximation of fairness and efficiency. This contrasts starkly with the findings in offline settings, where it is possible to efficiently obtain an allocation satisfying EF1, MMS, or min-USC [18].
@@
(Lemma 2 in [18]) A cost function $c$ is supermodular and binary if and only if the function $g$, $g(S)=|S|-c(S)$ for any set $S$, is a matroid rank function.
@@
For supermodular binary cost functions, no deterministic online algorithms can achieve $\alpha$-EF1 and completeness for any $\alpha\geq 1$.
@@
Consider an online allocation instance, there are two agents $\{1,2\}$ with supermodular binary cost functions $c_{1}$ and $c_{2}$. Based on the relationship between the supermodular and binary functions and the mathematical rank function in Lemma 10, we construct cost functions for the instance by constructing a matroid rank function. For $i\in\{1,2\}$, the online items in $T$ belongs to several categories $\{C^{i}_{1},\cdots,C^{i}_{l}\}$. In Figure 3, categories are distinguished by colours, and edges of different colours represent items of different categories for adjacent agents. For agent $i\in\{1,2\}$, consider the corresponding partition matroid $\mathcal{M}_{i}=\{T,\mathcal{F}_{i}\}$, $\mathcal{F}_{i}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~{}k\in[l]\}$. Let the function $r_{i}$ be the rank function of matroid $\mathcal{M}_{i}$, the cost function $c_{i}(\cdot)$ is defined as $c_{i}(S)=|S|-r_{i}(S)$ for any item set $S\in T$. For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are 00. W.l.o.g., we allocate item $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 00. We will discuss the following two cases.
@@
Case 1: we allocate $e_{2}$ to agent $1$. For the third item $e_{3}$, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00. Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases. Case 1-1: we allocate $e_{3}$ to agent $1$. When the fourth item $e_{4}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00. To not break the requirement of EF1, we must allocate item $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00. To avoid breaking the requirement of EF1, we must allocate item $e_{5}$ to agent $2$. When the sixth item $e_{6}$ arrives, the marginal utility for agent $1$ and agent $2$ is $1$. If we allocate $e_{6}$ to agent $1$, $A^{6}_{1}=\{e_{1},e_{2},e_{3},e_{6}\}$, $A^{6}_{2}=\{e_{4},e_{5}\}$, then we have $c_{1}(A^{6}_{2})=0<1=c_{1}(A^{6}_{1}\setminus\{e\})$. Else, we allocate $e_{6}$ to agent $2$, $A^{6}_{1}=\{e_{1},e_{2},e_{3}\}$, $A^{6}_{2}=\{e_{4},e_{5},e_{6}\}$, then we have $c_{2}(A^{6}_{1})=0<1=c_{2}(A^{6}_{2}\setminus\{e\})$. Case 1-2: we allocate $e_{3}$ to agent $2$. When the fourth item $e_{4}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00. Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases. Case 1-2-1: we allocate $e_{4}$ to agent $1$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{5}$ to agent $2$. When the sixth item $e_{6}$ arrives, the marginal utility for agent $1$ and agent $2$ is $1$. If we allocate $e_{6}$ to agent $1$, $A^{6}_{1}=\{e_{1},e_{2},e_{4},e_{6}\}$, $A^{6}_{2}=\{e_{3},e_{5}\}$, then we have $c_{1}(A^{6}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$ for any $e\in A^{5}_{1}$. Else, we allocate $e_{6}$ to agent $2$, $A^{6}_{1}=\{e_{1},e_{2},e_{4}\}$, $A^{6}_{2}=\{e_{3},e_{5},e_{6}\}$, then we have $c_{2}(A^{6}_{1})=0<1=c_{2}(A^{6}_{2}\setminus\{e\})$ for any $e\in A^{6}_{2}$. Case 1-2-2: we allocate $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{5}$ to agent $1$. When the sixth item $e_{6}$ arrives, the marginal utility for agent $1$ and agent $2$ is $1$. If we allocate $e_{6}$ to agent $1$, $A^{6}_{1}=\{e_{1},e_{2},e_{5},e_{6}\}$, $A^{6}_{2}=\{e_{3},e_{4}\}$, then we have $c_{1}(A^{6}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$ for any $e\in A^{5}_{1}$. Else, we allocate $e_{6}$ to agent $2$, $A^{6}_{1}=\{e_{1},e_{2},e_{5}\}$, $A^{6}_{2}=\{e_{3},e_{4},e_{6}\}$, then we have $c_{2}(A^{6}_{1})=0<1=c_{2}(A^{6}_{2}\setminus\{e\})$ for any $e\in A^{6}_{2}$.
@@
Case 2: we allocate $e_{2}$ to agent $2$. For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are 00. Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases. Case 2-1: we allocate $e_{3}$ to agent $1$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. If we allocate $e_{5}$ to agent $1$, $A^{5}_{1}=\{e_{1},e_{3},e_{5}\}$, $A^{5}_{2}=\{e_{2},e_{4}\}$, then we have $c_{1}(A^{5}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$. Else, we allocate $e_{5}$ to agent $2$, $A^{5}_{1}=\{e_{1},e_{3}\}$, $A^{5}_{2}=\{e_{2},e_{4},e_{5}\}$, then we have $c_{2}(A^{5}_{1})=0<1=c_{2}(A^{5}_{2}\setminus\{e\})$. Case 2-2: we allocate $e_{3}$ to agent $2$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{4}$ to agent $1$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. If we allocate $e_{5}$ to agent $1$, $A^{5}_{1}=\{e_{1},e_{4},e_{5}\}$, $A^{5}_{2}=\{e_{2},e_{3}\}$, then we have $c_{1}(A^{5}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$. Else, we allocate $e_{5}$ to agent $2$, $A^{5}_{1}=\{e_{1},e_{4}\}$, $A^{5}_{2}=\{e_{2},e_{3},e_{5}\}$, then we have $c_{2}(A^{5}_{1})=0<1=c_{2}(A^{5}_{2}\setminus\{e\})$. ∎
Theorem 11.
For supermodular binary cost functions, no deterministic online algorithms can achieve $\alpha$-MMS or $\alpha$-min-USC, and completeness for any $\alpha\geq 1$.
Proof.
@@
Consider an online allocation instance, there are two agents $\{1,2\}$ with supermodular binary cost functions $c_{1},c_{2}$. For $i\in\{1,2\}$, the online items in $T$ belongs to several categories $C^{i}_{1},\cdots,C^{i}_{l}$. In Figure 2, categories are distinguished by colours, and edges of different colours represent items of different categories for adjacent agents. For agent $i\in\{1,2\}$, consider the corresponding partition matroid $\mathcal{M}_{i}=\{T,\mathcal{F}_{i}\}$, $\mathcal{F}_{i}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~{}k\in[l]\}$. Let the function $r_{i}$ be the rank function of matroid $\mathcal{M}_{i}$, the cost function $c_{i}(\cdot)$ is defined as $c_{i}(S)=|S|-r_{i}(S)$ for any item set $S\in T$. For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are 00. We allocate item $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 00. We will discuss the following two cases.
@@
Case 1: We allocate $e_{2}$ to agent $1$. For the third item $e_{3}$, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00. In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{3}$ to agent $2$. When the fourth item $e_{4}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. No matter which agent is assigned, the cost of its allocation is $1$, but $\mathsf{MMS}_{1}(T)=\mathsf{MMS}_{2}(T)=0$, $\min USC(T)=0$.
@@
Case 2: We allocate $e_{2}$ to agent $2$. For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are 00. Continue to discuss in two sub-cases. Case 2-1: We allocate $e_{3}$ to agent $2$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are $1$. No matter which agent is assigned, the cost of its allocation is $1$, but $\mathsf{MMS}_{1}(T)=\mathsf{MMS}_{2}(T)=0$, $\min USC(T)=0$. Case 2-2: We allocate $e_{3}$ to agent $1$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are 00. In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 00. In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{5}$ to agent $2$. When the sixth item $e_{6}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. No matter which agent is assigned, the cost of its allocation is $1$, but $\mathsf{MMS}_{1}(T)=\mathsf{MMS}_{2}(T)=0$, $\min USC(T)=0$. ∎
4.3 Additive Bi-valued Costs

### v2

@@ 4.2.1 Additive Binary Costs
Output: An complete EF1 allocation $\bf A$}
1 Initialize: let $\pi=(\pi_{1},\ldots,\pi_{n})$ be an arbitrary order of $n$ agents;
2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
@@ 14 $\pi=(\pi_{2},\ldots,\pi_{i-1},\pi_{i+1},\ldots,\pi_{n},\pi_{1})$;
Given an online fair allocation instance $I=(T,N,(c_{i})_{i\in N})$, where $c_{i}$ is additive and binary for all $i\in N$. Fix an arbitrary round $k$, let $T=\{e_{1},\ldots,e_{k}\}$ be the set of items that have already arrived, denoted by ${\bf A^{k}}=(A^{k}_{1},\ldots,A^{k}_{n})$ the allocation at the end of round $k$. For every item $e\in T$, if there exists one agent such that the cost about $e$ of some agent is 0, the algorithm must allocate $e$ to some agent $a$ with $v_{a}(e)=0$. Otherwise, the costs of all agents are $1$. Regardless of how it is allocated, the increment brought to utilitarian social welfare is $1$, so $A^{k}$ is a min-USC allocation.
@@ 4.2.2 Supermodular Binary Costs
For supermodular binary cost functions, where for any agent $i\in N$, $c_{i}$ is supermodular and the marginal cost of any item is 0 or 1, we first obtain a significant negative result regarding the approximation of fairness and efficiency. This contrasts starkly with the findings in offline settings, where it is possible to efficiently obtain an allocation satisfying EF1, MMS, or min-USC Barman et al. (2023).
@@
(Lemma 2 in Barman et al. (2023)) A cost function $c$ is supermodular and binary if and only if the function $g$, g(S)=|S|c(S)g(S)=|S|-c(S) for any set $S$, is a matroid rank function.
@@
For any $\alpha\geq 1$, with supermodular binary cost functions, no deterministic online algorithms can achieve
  • $\alpha$-EF1 and completeness,
  • or $\alpha$-MMS and completeness,
  • or $\alpha$-min-USC and completeness.
@@
First, let us consider EF1. Consider an online allocation instance; there are two agents $\{1,2\}$ with supermodular binary cost functions $c_{1}$ and $c_{2}$.
@@
Based on the relationship between the supermodular and binary functions and the mathematical rank function in Lemma 7, we construct cost functions for the instance by constructing a matroid rank function. For $i\in\{1,2\}$, the online items in $T$ belongs to several categories $\{C^{i}_{1},\ldots,C^{i}_{l}\}$. In Figure 3, categories are distinguished by colours, and edges of different colours represent items of different categories for adjacent agents. For agent $i\in\{1,2\}$, consider the corresponding partition matroid $\mathcal{M}_{i}=\{T,\mathcal{F}_{i}\}$, $\mathcal{F}_{i}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~k\in[l]\}$. Let the function $r_{i}$ be the rank function of matroid $\mathcal{M}_{i}$, the cost function $c_{i}(\cdot)$ is defined as $c_{i}(S)=|S|-r_{i}(S)$ for any item set $S\in T$. For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are 0. W.l.o.g., we allocate item $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 0. We will discuss the following two cases.
@@
Case 1: we allocate $e_{2}$ to agent $1$. For the third item $e_{3}$, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0. Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases. Case 1-1: we allocate $e_{3}$ to agent $1$. When the fourth item $e_{4}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0. To not break the requirement of EF1, we must allocate item $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0. To avoid breaking the requirement of EF1, we must allocate item $e_{5}$ to agent $2$. When the sixth item $e_{6}$ arrives, the marginal utility for agent $1$ and agent $2$ is $1$. If we allocate $e_{6}$ to agent $1$, $A^{6}_{1}=\{e_{1},e_{2},e_{3},e_{6}\}$, $A^{6}_{2}=\{e_{4},e_{5}\}$, then we have $c_{1}(A^{6}_{2})=0<1=c_{1}(A^{6}_{1}\setminus\{e\})$. Else, we allocate $e_{6}$ to agent $2$, $A^{6}_{1}=\{e_{1},e_{2},e_{3}\}$, $A^{6}_{2}=\{e_{4},e_{5},e_{6}\}$, then we have $c_{2}(A^{6}_{1})=0<1=c_{2}(A^{6}_{2}\setminus\{e\})$. Case 1-2: we allocate $e_{3}$ to agent $2$. When the fourth item $e_{4}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0. Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases. Case 1-2-1: we allocate $e_{4}$ to agent $1$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{5}$ to agent $2$. When the sixth item $e_{6}$ arrives, the marginal utility for agent $1$ and agent $2$ is $1$. If we allocate $e_{6}$ to agent $1$, $A^{6}_{1}=\{e_{1},e_{2},e_{4},e_{6}\}$, $A^{6}_{2}=\{e_{3},e_{5}\}$, then we have $c_{1}(A^{6}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$ for any $e\in A^{5}_{1}$. Else, we allocate $e_{6}$ to agent $2$, $A^{6}_{1}=\{e_{1},e_{2},e_{4}\}$, $A^{6}_{2}=\{e_{3},e_{5},e_{6}\}$, then we have $c_{2}(A^{6}_{1})=0<1=c_{2}(A^{6}_{2}\setminus\{e\})$ for any $e\in A^{6}_{2}$. Case 1-2-2: we allocate $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{5}$ to agent $1$. When the sixth item $e_{6}$ arrives, the marginal utility for agent $1$ and agent $2$ is $1$. If we allocate $e_{6}$ to agent $1$, $A^{6}_{1}=\{e_{1},e_{2},e_{5},e_{6}\}$, $A^{6}_{2}=\{e_{3},e_{4}\}$, then we have $c_{1}(A^{6}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$ for any $e\in A^{5}_{1}$. Else, we allocate $e_{6}$ to agent $2$, $A^{6}_{1}=\{e_{1},e_{2},e_{5}\}$, $A^{6}_{2}=\{e_{3},e_{4},e_{6}\}$, then we have $c_{2}(A^{6}_{1})=0<1=c_{2}(A^{6}_{2}\setminus\{e\})$ for any $e\in A^{6}_{2}$.
@@
Case 2: we allocate $e_{2}$ to agent $2$. For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are 0. Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases. Case 2-1: we allocate $e_{3}$ to agent $1$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. If we allocate $e_{5}$ to agent $1$, $A^{5}_{1}=\{e_{1},e_{3},e_{5}\}$, $A^{5}_{2}=\{e_{2},e_{4}\}$, then we have $c_{1}(A^{5}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$. Else, we allocate $e_{5}$ to agent $2$, $A^{5}_{1}=\{e_{1},e_{3}\}$, $A^{5}_{2}=\{e_{2},e_{4},e_{5}\}$, then we have $c_{2}(A^{5}_{1})=0<1=c_{2}(A^{5}_{2}\setminus\{e\})$. Case 2-2: we allocate $e_{3}$ to agent $2$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{4}$ to agent $1$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. If we allocate $e_{5}$ to agent $1$, $A^{5}_{1}=\{e_{1},e_{4},e_{5}\}$, $A^{5}_{2}=\{e_{2},e_{3}\}$, then we have $c_{1}(A^{5}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$. Else, we allocate $e_{5}$ to agent $2$, $A^{5}_{1}=\{e_{1},e_{4}\}$, $A^{5}_{2}=\{e_{2},e_{3},e_{5}\}$, then we have $c_{2}(A^{5}_{1})=0<1=c_{2}(A^{5}_{2}\setminus\{e\})$.
@@
Next, let us consider MMS or USC. Consider an online allocation instance; there are two agents $\{1,2\}$ with supermodular binary cost functions $v_{1}v_{2}$. For $i\in\{1,2\}$, the online items in $T$ belongs to several categories $C^{i}_{1},\ldots,C^{i}_{l}$. In Figure 4, categories are distinguished by colours, and edges of different colours represent items of different categories for adjacent agents. For agent $i\in\{1,2\}$, consider the corresponding partition matroid $\mathcal{M}_{i}=\{T,\mathcal{F}_{i}\}$, $\mathcal{F}_{i}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~k\in[l]\}$. Let the function $r_{i}$ be the rank function of matroid $\mathcal{M}_{i}$, the cost function $c_{i}(\cdot)$ is defined as $c_{i}(S)=|S|-r_{i}(S)$ for any item set $S\in T$. For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are 0. We allocate item $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 0. We will discuss the following two cases.
@@
Case 1: we allocate $e_{2}$ to agent $1$. For the third item $e_{3}$, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0. In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{3}$ to agent $2$. When the fourth item $e_{4}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. No matter which agent is assigned, the cost of its allocation is $1$, but $\mathsf{MMS}_{1}(T)=\mathsf{MMS}_{2}(T)=0$, $\min USC(T)=0$.
@@
Case 2: we allocate $e_{2}$ to agent $2$. For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are 0. Continue to discuss in two sub-cases Case 2-1: we allocate $e_{3}$ to agent $2$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are $1$. No matter which agent is assigned, the cost of its allocation is $1$, but $\mathsf{MMS}_{1}(T)=\mathsf{MMS}_{2}(T)=0$, $\min USC(T)=0$. Case 2-2: we allocate $e_{3}$ to agent $1$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are 0. In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 0. In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{5}$ to agent $2$. When the sixth item $e_{6}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. No matter which agent is assigned, the cost of its allocation is $1$, but $\mathsf{MMS}_{1}(T)=\mathsf{MMS}_{2}(T)=0$, $\min USC(T)=0$. ∎
4.3 Additive Personalized Bi-valued Costs